In a question, an inclined surface is inclined at arcsin 1/3 to the horizontal surface. What does this arcsin 1/3 means? 

Grade:12

4 Answers

AskIITians Expert PRASAD IIT Kharagpur
14 years ago

 

arcsin = angle

Let, @ = arcsin(1/3)

@ = angle between hypotenuse and the horizontal surface.

Then sin @ = 1/3

As sin @ = opposite side / hypotenuse

we have hypotenuse is thrice the length of the side opposite angle @.

Ramesh V
14 years ago


  An abbreviation for inverse sine.
 arsin(x) or arcsin(x) means the inverse sine of x, i.e. the number whose sine is x.
 Occasionally asin(x) is used to mean the same thing.

 

here its sin-1(1/3)

AskIITians Expert Hari Shankar IITD
14 years ago

arcsin is sin inverse.

theta = arcsin(1/3) means sin(theta) = 1/3

Sameer Satyam
14 years ago

Just another way to write sine inverse.......